OpenStudy (steve816):

When you square root an inequality, such as this one: \[x^2 \le 4\] Do you flip the inequality sign? How would I go about solving this?

OpenStudy (sshayer):

\[\left| x \right|\leq 2\] \[-2 \leq x \leq 2\]

OpenStudy (sshayer):

\[x^2 \leq a^2\] \[\left| x \right| \leq a,-a \leq x \leq a\] \[x^2 \geq a^2,x \leq -a,x \geq a\]

OpenStudy (steve816):

How come I never learned about this??? So is this a rule that you have to memorize, because logically, it isn't making sense to me

OpenStudy (sshayer):

i tell you another way. \[x^2 \leq 4,-2 \leq x \leq 2\] or x lies between -2 and 2 take any number between -2 and 2 say-1 ,0 ,1 \[\left( -1 \right)^2=1 \leq 4\] \[0^2=0 \leq 4\] \[1^2=1 \leq4\] now i take any number beyond this interval say -3 \[\left( -3 \right)^2=9 \geq 4\] similarly \[3^2=9 \geq 4\] now it should be clear to you.

OpenStudy (sshayer):

i have given you general formula.

jimthompson5910 (jim_thompson5910):

Naturally it might be tempting to say \[\Large \sqrt{x^2} = x\] which is not true for all real numbers. It's only true if x is not negative. To ensure the output is nonnegative, we use absolute value brackets So it's really \[\Large \sqrt{x^2} = |x|\] which is true for all real values x (even when x is negative)

jimthompson5910 (jim_thompson5910):

So that's why after square rooting both sides, we go from \[\Large x^2 < 4\] to \[\Large |x| < 2\]

jimthompson5910 (jim_thompson5910):

Then you'll use these rules http://www.mhhe.com/math/devmath/dugopolski/inter/instructor/ppt/ppt02/img017.jpg to break things down further

OpenStudy (kainui):

Just to show some concrete examples of when this is true for integers (also true for nonintegers like 1/2 too): \[x^2 \le 4\] I'll go ahead and plug in all of the integers: \[(-2)^2\le 4\]\[(-1)^2 \le 4\]\[0^2 \le 4\]\[1^2\le 4\]\[2^2\le 4\] we can see since the square root makes negative numbers positive, it sorta "bounces" or "folds" the boundary of the interval [-2,2] back onto itself [0,2] So when you take the square root of all the values to keep the interval the same since\(\sqrt{4}=2\)and \(\sqrt{4}=-2\), we have: \[-2 \le -2 \le 2\]\[-2 \le -1 \le 2\]\[-2 \le 0 \le 2\]\[-2 \le 1 \le 2\]\[-2 \le 2 \le 2\] This might confuse you at first, but I think it should help make it make sense for you @steve816 if you give it a little time and/or ask me questions I'll be happy to help.

OpenStudy (steve816):

Thanks for the response @kainui, it is much more clear now. I have no idea why I didn't know about this until I started calculus, and seems like most teachers never go over this case during inequality unit in algebra.